OpenStudy (anonymous):

Use substitution to solve the system 5x + 8y= −25 4x − 5y= −20 (x,y)= ( , )

OpenStudy (saifoo.khan):

we can also use substitution. :P

OpenStudy (anonymous):

Hero beat me to it... touche

OpenStudy (anonymous):

by the looks of it, waaay before i did.. good job

hero (hero):

1. Solve for y: y = (-25 - 5x)/8 y = (20 + 4x)/5 2. Substitute y = y: (-25 - 5x)/8 = (20 + 4x)/5 3. Cross Multiply: 5(-25 - 5x) = 8(20+4x) 4. Continue Solving for x: -125 - 25x = 160 + 32x 57x = -285 x = -285/57 = -5 y = 0

hero (hero):

Robtobey always tries to come in behind me. I made a mistake, but corrected it

hero (hero):

Did you post something?

OpenStudy (anonymous):

me? i was about to, im just looking it over trying it myself,but be prepared if i end up asking u a question later bout it

hero (hero):

Okay. It's not the traditional version of substitution, but it's still substitution. It can be seen as substitution or elimination really because you're substituting the equation of y in terms of x and eliminating the y variable at the same time.

hero (hero):

So in a way, it's the combo version

OpenStudy (anonymous):

yeah it did look different than what im use to

hero (hero):

It's just solving for y and setting it equal to each other. The reason I do it that way is sometimes when you use traditional elimination or substitution methods, you can run into a trap that give you false answers. Using my method, you don't run into such traps and you can easily see if there is a solution or if there is no solution.

OpenStudy (anonymous):

will it work for most problems like that?

hero (hero):

It will work for all of them with no problem. The only thing is.... If you solve for y and the denominators are the same, when you cross multiply, be sure to properly cancel, otherwise you will get a wrong answer